# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 6.5 Application domains
In this Table 5, we link the discussed methodologies and applications to their respective domains. It is evident that the medical domain, followed by the automotive field, is the most prominent for probabilistic segmentation tasks. The high-stakes nature of both domains contributes to their prominence, as uncertainty-aware models are especially vital in safety-critical applications. Specifically, the delineation of lung nodules in CT images are ambiguous. The LIDC-IDRI dataset is a multi-annotated Lung CT dataset, which serves as a good benchmark for quantifying observer variability and is reflected in the table.
@@
Outdoor scenes Kahl et al. (2024); Kohl et al. (2019b); Kassapis et al. (2021); Gao et al. (2022) Sinha et al. (2019); Kahl et al. (2024); Kasarla et al. (2019); Kim et al. (2021); Xie et al. (2022); Wu et al. (2021) Pavlitskaya et al. (2020); Hasan & Linte (2022b); Sander et al. (2019); Ng et al. (2022); Hasan & Linte (2022a); Mehrtash et al. (2020); Kendall et al. (2016); Kahl et al. (2024); Holder & Shafique (2021); Qi et al. (2023) Ng et al. (2022); Huang et al. (2018); Zbinden et al. (2023); Mukhoti & Gal (2018); Amit et al. (2021); Hoogeboom et al. (2021); Valada et al. (2017); Rakic et al. (2024)
@@
Remote sensing García Rodríguez et al. (2020) Dechesne et al. (2021; 2021); Kampffmeyer et al. (2016); Dechesne et al. (2021)
Microscopy Kohl et al. (2019a); Zepf et al. (2023a); Schmidt et al. (2023); Philps et al. (2024) Khalili et al. (2024); Li & Alstrøm (2020); Yang et al. (2017); Sadafi et al. (2019) Linmans et al. (2020); Iwamoto et al. (2021) Bogensperger et al. (2023); Amit et al. (2021)
Brain MRI Philps et al. (2024); Bhat et al. (2023); Hu et al. (2023); Philps et al. (2024); Savadikar et al. (2021); Zhang et al. (2022c) Zepf et al. (2023b); Ma et al. (2024b); Shen et al. (2021) Roy et al. (2018b); Mehrtash et al. (2020); Jungo & Reyes (2019); Eaton-Rosen et al. (2018); Wang et al. (2019b); Jungo et al. (2020); Whitbread & Jenkinson (2022); Monteiro et al. (2020); Lambert et al. (2022); Jungo et al. (2018b; a) Wolleb et al. (2021); Wu et al. (2023b; a); Chen et al. (2023a); Wang et al. (2019a); Kamnitsas et al. (2018); Larrazabal et al. (2021b); Wang et al. (2019b); Whitbread & Jenkinson (2022)
@@
Dermoscopy Li & Alstrøm (2020); Gorriz et al. (2017) Zepf et al. (2023b); Jungo & Reyes (2019); DeVries & Taylor (2018); Czolbe et al. (2021) Wu et al. (2023a); Zepf et al. (2023a)
Lung CT Rafael-Palou et al. (2021); Gao et al. (2022); Valiuddin et al. (2024b); Chen et al. (2023a); Zhang et al. (2022b); Monteiro et al. (2020); Qiu & Lui (2020); Kohl et al. (2019a); Hu et al. (2022); Zhang et al. (2022c); Gantenbein et al. (2020); Viviers et al. (2023b; c); Valiuddin et al. (2024a); Zbinden et al. (2023); Zepf et al. (2024); Kahl et al. (2024); Bhat et al. (2022a); Long et al. (2021b); Valiuddin et al. (2021); Rahman et al. (2023); Kohl et al. (2019b); Hu et al. (2023); Bhat et al. (2023); Kassapis et al. (2021); Long et al. (2021a); Selvan et al. (2020); Baumgartner et al. (2019) Kahl et al. (2024); Hoebel et al. (2020) Kahl et al. (2024); Czolbe et al. (2021); Chotzoglou & Kainz (2019) Hu et al. (2022); Rakic et al. (2024); Zhang et al. (2022a)
Various CT Valiuddin et al. (2024b); Hu et al. (2023; 2023) Hiasa et al. (2019) Hoebel et al. (2019); LaBonte et al. (2019; 2019) Viviers et al. (2023a)
@@ 7 Discussion
This section builds on the reviewed theory and literature to guide researchers in selecting methods suited to their goals. Furthermore, it highlights key gaps and limitations, offering insights for addressing current challenges in the field and identifying promising directions for future research. Our main discussion points can be summarized as follows:
  • The interpretation and quantification of epistemic and aleatoric uncertainty remain complex and warrant careful reconsideration, despite the literature often indicating a clear distinction between them. (Section 7.1).
  • Spatial uncertainty introduces additional challenges, including the accurate estimation of structural entropy, the aggregation of pixel-wise uncertainties, and increased complexity in volumetric segmentation tasks (Section 7.2).
  • Lack of standardization across the downstream tasks remains a major issue in the field, hindering fair benchmarking and limiting further progress (Section 7.3).
  • Active Learning is a promising application of uncertainty quantification, reducing annotation costs while maintaining or even improving model generalization (Section 7.4).
  • Segmentation backbones commonly used in the field tend to be outdated, with Transformer-based architectures having surpassed them in performance yet remaining underutilized in probabilistic dense predictions (Section 7.5).
  • Aleatoric uncertainty is often modeled using conditional generative approaches; as this area advances rapidly, its developments should be promptly integrated into probabilistic segmentation research (Section 7.6).
  • Alternatives to epistemic uncertainty modeling warrant further exploration due to longstanding criticisms of the predominantly used Monte Carlo dropout, questioning its ability to capture true model uncertainty (Section 7.7).
@@ 7.1 Disentangling uncertainties
Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric). Importantly, this distinction relates to the suggestion whether the uncertainty is reducible or not. Furthermore, Equation( 2) indicates this can be achieved by introducing stochasticity into the model parameters or the features. Seemingly straightforward, sufficient literature and discussion in the scientific community exist related to the applicability of this taxonomy in practical cases. As paraphrased from Kirchhof et al. (2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”. We further explore this concept and show how confusion and contradiction have led to common pitfalls in related studies. To address these issues, we highlight key nuances in uncertainty modeling and contextualize this to segmentation problems by providing recommendations for future research.
7.1.1 Nuances in uncertainty modeling
Interpretation of uncertainty is task dependent
@@ 7.1.1 Nuances in uncertainty modeling
It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context. For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Sections 6.3 and 6.4). Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2). Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty Depeweg et al. (2018), such as commonly done in Active Learning (Section 6.2). Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk Apostolakis (1990); Helton (1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1). However, it was highlighted that the decision to model a specific type of uncertainty was largely influenced by how the problem was framed. To avoid confusion, this task-dependency should be noted when interpreting uncertainty.
Interpretation of uncertainty is model dependent
Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian & Ditlevsen, 2009). For instance, the expected entropy w.r.t. a stochastic parameter variable should represent the aleatoric uncertainty (Equation (3)). However, when employing generative models, this behavior changes. The latent (feature) stochasticity variable is employed to model the aleatoric uncertainty. Hence, the remaining uncertainty, i.e. epistemic, is encapsulated in the expected entropy w.r.t. this latent variable (Equation (13)). Hence, the calibration network of the GAN-based CAR model of Kassapis et al. (2021) (discussed in Section 4.2.1) actually quantified the epistemic rather than aleatoric uncertainty. This is dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA (Wang et al., 2019a; Ayhan & Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Kahl et al., 2024; Hu et al., 2022; Whitbread & Jenkinson, 2022; Roshanzamir et al., 2023). The key message thus far is that, after selecting the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty type.
Uncertainties are intertwined
@@
Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion. A systematic study has found that the ability to separate the uncertainties can strongly depend on the data Kahl et al. (2024). These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods de Jong et al. (2024); Mucsányi et al. (2024), indicating that they represent uncertainty of similar kind. Kendall & Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other. In probabilistic models, we can also find that the epistemic encapsulation methods have been used to model observer variability. Furthermore, using a combination of aleatoric end epistemic modeling techniques can also beneficial (Gao et al., 2022). These findings further highlight the notion of the uncertainties being intertwined. On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions Kingma et al. (2015), further supporting this argument.
Encapsulating vs. quantifying
@@
A common misconception pertains that modeling a specific uncertainty also implies that the predicted output represents that same uncertainty. This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e. targeting an intended uncertainty type, versus quantifying, i.e. measuring the modeled uncertainty in isolation. For example, the estimation of mutual information is required to express the epistemic uncertainty in isolation. Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties (Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018). At the very most, the output correlates with an increase in epistemic uncertainty. However, given the entanglement of both uncertainties, part of the predictive variance can be accredited to noise in the data. This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance Zhang et al. (2022c); Monteiro et al. (2020); Valiuddin et al. (2024b); Gao et al. (2022). However, epistemic uncertainty is often assumed to be explained away or assumed to be comparatively negligible with large datasets. Finally, we have found it to be common to interchangeably use the total and model uncertainty, especially in the context of Monte Carlo Dropout methods Kendall et al. (2016); Roy et al. (2018b); DeVries & Taylor (2018); Mehrtash et al. (2020); Burmeister et al. (2022); Iwamoto et al. (2021). To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
7.1.2 Pragmatic use of uncertainty
Unless being the core contribution, there seems to be no compelling need to investigate uncertainty disentanglement for the applications of uncertainty in dense prediction. If performance on the downstream application is improved, it can be assumed that modeled uncertainty correlates with changes in variance of the predictive distribution. In such cases, no claims need to be made on the uncertainty type and neither is it required to isolate the two. Hence, as described in previous work, aleatoric and epistemic uncertainty should be used as to communicate whether the uncertainty is reducible or not (Der Kiureghian & Ditlevsen, 2009; Faber, 2005), to then select appropriate modeling methods.
As argued by Kahl et al. (2024), claiming to quantify an uncertainty type would ideally require validation across all relevant applications. For example, validation of epistemic uncertainty requires a clear distribution shift in the data (domain, semantic, covariate etc.), but it is often found in literature that either no distribution shift or the required metrics are used (Zhang et al., 2022a; Wang et al., 2019a; Postels et al., 2019; Mukhoti & Gal, 2018; Mobiny et al., 2021; Whitbread & Jenkinson, 2022). As discussed in Section 6.1, aleatoric uncertainty should be validated with multiple labels and appropriate metrics, but can often found to be evaluated on a single annotator or without the use of distribution-level statistics Wang et al. (2019a); Whitbread & Jenkinson (2022); Kendall & Gal (2017); Liu et al. (2022b); Schmidt et al. (2023); Savadikar et al. (2021). Hence, simply avoiding any serious claims related to either type alleviates this burden, and allows one to focus solely on the task at hand. In contrast to the work of Kahl et al. (2024), however, we do not consider calibration of models itself as a downstream task, but rather as a tool to gauge model reliability to be later used for other tasks. As a consequence, the pixel-wise uncertainties can be later used for possible downstream applications.

### v8

@@ 6.5 Application domains
In Table 5, we link the discussed methodologies and tasks to their respective application domains. It is evident that the medical domain, followed by the automotive field, is the most prominent for probabilistic segmentation tasks. The high-stakes nature of both domains contributes to their prominence, as uncertainty-aware models are especially vital in safety-critical applications. Several benchmark datasets are central to these applications. For instance, the LIDC-IDRI dataset (Armato, 2011), with its multiple expert delineations of lung nodules in CT scans, is a key benchmark for quantifying observer variability. In the autonomous driving domain, outdoor scene understanding applications widely utilize the Cityscapes dataset. Similarly, MRI-based research is often advanced by datasets like BraTS or challenges such as QUBIQ (Menze et al., 2021).
@@
Outdoor scenes Kahl et al. (2024); Kohl et al. (2019b); Kassapis et al. (2021); Gao et al. (2022) Sinha et al. (2019); Kahl et al. (2024); Kasarla et al. (2019); Kim et al. (2021); Xie et al. (2022); Wu et al. (2021) Pavlitskaya et al. (2020); Hasan and Linte (2022b); Sander et al. (2019); Ng et al. (2022); Hasan and Linte (2022a); Mehrtash et al. (2020); Kendall et al. (2016); Kahl et al. (2024); Holder and Shafique (2021); Qi et al. (2023) Ng et al. (2022); Huang et al. (2018); Zbinden et al. (2023); Mukhoti and Gal (2018); Amit et al. (2021); Hoogeboom et al. (2021); Valada et al. (2017); Rakic et al. (2024)
@@
Remote sensing García Rodríguez et al. (2020) Dechesne et al. (2021); Kampffmeyer et al. (2016); Dechesne et al. (2021)
Microscopy Kohl et al. (2019a); Zepf et al. (2023a); Schmidt et al. (2023); Philps et al. (2024) Khalili et al. (2024); Li and Alstrøm (2020); Yang et al. (2017); Sadafi et al. (2019) Linmans et al. (2020); Iwamoto et al. (2021) Bogensperger et al. (2023); Amit et al. (2021)
Brain MRI Philps et al. (2024); Bhat et al. (2023); Hu et al. (2023); Philps et al. (2024); Savadikar et al. (2021); Zhang et al. (2022d) Zepf et al. (2023b); Ma et al. (2024b); Shen et al. (2021) Roy et al. (2018b); Mehrtash et al. (2020); Jungo and Reyes (2019); Eaton-Rosen et al. (2018); Wang et al. (2019b); Jungo et al. (2020); Whitbread and Jenkinson (2022); Monteiro et al. (2020); Lambert et al. (2022); Jungo et al. (2018b; a) Wolleb et al. (2021); Wu et al. (2023b; a); Chen et al. (2023a); Wang et al. (2019a); Kamnitsas et al. (2018); Larrazabal et al. (2021a); Wang et al. (2019b); Whitbread and Jenkinson (2022)
@@
Dermoscopy Li and Alstrøm (2020); Gorriz et al. (2017) Zepf et al. (2023b); Jungo and Reyes (2019); DeVries and Taylor (2018); Czolbe et al. (2021) Wu et al. (2023a); Zepf et al. (2023a)
Lung CT Rafael-Palou et al. (2021); Gao et al. (2022); Valiuddin et al. (2024b); Chen et al. (2023a); Zhang et al. (2022c); Monteiro et al. (2020); Qiu and Lui (2020); Kohl et al. (2019a); Hu et al. (2022); Zhang et al. (2022d); Gantenbein et al. (2020); Viviers et al. (2023c; a); Valiuddin et al. (2024a); Zbinden et al. (2023); Zepf et al. (2024); Kahl et al. (2024); Bhat et al. (2022a); Long et al. (2021a); Valiuddin et al. (2021); Rahman et al. (2023); Kohl et al. (2019b); Hu et al. (2023); Bhat et al. (2023); Kassapis et al. (2021); Long et al. (2021b); Selvan et al. (2020); Baumgartner et al. (2019) Kahl et al. (2024); Hoebel et al. (2020) Kahl et al. (2024); Czolbe et al. (2021); Chotzoglou and Kainz (2019) Hu et al. (2022); Rakic et al. (2024); Zhang et al. (2022a)
Various CT Valiuddin et al. (2024b); Hu et al. (2023) Hiasa et al. (2019) Hoebel et al. (2019); LaBonte et al. (2019) Viviers et al. (2023b)
@@ 7 Discussion
This section builds on the reviewed theory and literature to guide researchers in selecting methods suited to their goals. Furthermore, it highlights key gaps and limitations, offering insights to address current challenges and identify promising directions for future research.
Main discussion points: The interpretation of epistemic and aleatoric uncertainty remain complex and warrant careful reconsideration, despite the literature often suggesting a clear distinction (Section 7.1). Challenges related to spatial coherence remain unaddressed (Section 7.2). The application-driven field lacks standardization across the downstream tasks, hindering benchmarking and further progress (Section 7.3). Uncertainty-based Active Learning is promising and has many open challenges (Section 7.4). There exist various limitations and trade-offs across methods (Section 7.5). Understanding the critical influence of data characteristics on uncertainty models remains a relatively under-explored area of research (Section 7.6). The field of segmentation could advance significantly by treating the backbone as a critical modeling choice and embracing contemporary architectures (Section 7.7). The segmentation community should actively explore and integrate novel uncertainty quantification methodologies (Section 7.8). Uncertainty modeling should address the growing importance of multiclass, instance, and panoptic segmentation problems (Section 7.9).
@@ 7.1 Disentangling uncertainties
Proper disentanglement of the uncertainties allows to accredit high-entropy outputs to either model ignorance about the data (epistemic), or inherent noise in the data generation process (aleatoric). As Equation (2) implies, this can be achieved by introducing stochasticity into either the model’s parameters or its features. This fundamental distinction not only provides the organizing principle for our subsequent review but also aligns with the approaches empirically observed in the literature. Seemingly straightforward, sufficient literature and discussion in the scientific community exist related to the applicability of this taxonomy in practical cases. As paraphrased from Kirchhof et al. (April 28, 2025), “definitions of uncertainty resemble the shapes of clouds — clearly defined from a distance, but losing clarity during approach by dissolving in one another”. We further explore this, highlighting key nuances in uncertainty modeling and contextualize this to segmentation problems. To begin, we first establish the original definitions of these concepts to better understand why these concepts have lead to incorrect definitions, claims and general confusion.
Aleatoric uncertainty
reconsiders the non-deterministic relationship between $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, which implies that
with Dirac-delta function $\delta$, and mapping $F:\mathcal{X}\rightarrow\mathcal{Y}$. This relationship is characterized by the ambiguity in $\mathbf{X}$ and is inherently probabilistic, due to various reasons such as noise in the data (occlusions, sensor noise, insufficient resolution, etc.) or variability within a class (e.g. not all cats have tails). Hence, the observance of substantial aleatoric uncertainty can in some cases be inevitable, but may also signal the need for higher-quality data acquisition or shifting to another modality. The possible input dependency of the uncertainty develops into further categorization of either heteroscedastic (input dependent) or homoscedastic (input independent) aleatoric uncertainty.
Epistemic uncertainty
arises from model ignorance (i.e. related to the parameters). Consequently, epistemic uncertainty should not only be quantified but ideally also minimized. Epistemic uncertainty can be further categorized into two distinct types (Hüllermeier and Waegeman, 2021). The first type pertains uncertainty related to the capacity of the model. For example, under-parameterized models or approximate model posteriors, that can become too stringent to appropriately resemble the true posterior. The ambiguity on the best parameters given the limited capacity induces uncertainty in the learning process, which is also known as model uncertainty. Nevertheless, given the complexity of contemporary parameter-intensive CNNs, the model uncertainty is often assumed to be negligible. A more significant contribution to the epistemic uncertainty is due to the limited data availability, known as approximation uncertainty, and can often be reduced by collecting more data. Additionally, we have found it to be common to interchangeably use the total epistemic and solely model uncertainty, especially in the context of Monte Carlo Dropout methods (Kendall et al., 2016; Roy et al., 2018b; DeVries and Taylor, 2018; Mehrtash et al., 2020; Burmeister et al., 2022; Iwamoto et al., 2021). To prevent confusion in already established terminology, the term model uncertainty should be exclusively used for the induced ambiguity due to limited model complexity, rather than to denote general or the entire epistemic uncertainty.
Task dependency
@@
It is essential to consider that interpretation of the uncertainty source often differs depending on the modeling context. For example, epistemic uncertainty to be considered as the number of required models to fit the data (Wimmer et al., 2023), evident in ensembling and M-head approaches for tasks related to model introspection and generalization (Sections 6.3 and 6.4). Another perspective pertains model disagreement (Houlsby et al., 2011; Gal et al., 2017b; Kirsch, 2024), relating to the source of observer variability (Section 6.1) and methods that employ mixture-of-expert approaches (Section 5.1.2). Also, it has been simply considered to be the remainder after subtracting the aleatoric from the total uncertainty (Depeweg et al., 2018), such as commonly done in Active Learning (Section 6.2). Aleatoric uncertainty also knows several definitions such as being the Bayes-optimal residual risk (Apostolakis, 1990; Helton, 1997), but remains consistent in related literature as the inherent, irreducible noise in the data inducing variance in the ground-truth distribution (Section 6.1). To avoid confusion, this task-dependency should be considered when interpreting uncertainty.
Model dependency
Moreover, past literature as well as our overview indicate the influence of specific design choices on the nature of the modeled uncertainty (Der Kiureghian and Ditlevsen, 2009). For instance, the expected entropy w.r.t. a stochastic parameter variable is often assumed to represent the epistemic uncertainty as shown in Equation (3). However, when employing generative models, this behavior changes. As argued by Kahl et al. (2024), the predictive uncertainty can be decomposed as
In contrast to Equation (3), the mutual information term here encapsulates the aleatoric uncertainty, while the expected entropy reflects the epistemic uncertainty. This is because generative models operate under the assumption that the latent variable $\mathbf{Z}$ captures the inherent data ambiguity, thus representing the aleatoric uncertainty. The mutual information describes how much about this aleatoric uncertainty can be learned by observing the output $\mathbf{Y}$. If one were to know the optimal latent code $\mathbf{Z}$, the mutual information would be zero. In this case, all stochasticity in the output $\mathbf{Y}$ is perfectly represented by $\mathbf{Z}$, meaning the aleatoric uncertainty has been fully captured. Under this view, if all data uncertainty can be fully encapsulated by the latent variable $\mathbf{Z}$, there is no need to model this irreducible stochasticity in the predictive model $p(\mathbf{Y}|\mathbf{Z,\mathbf{x},\bm{\theta}})$, implying it corresponds to the epistemic uncertainty. This is consistent with the principle that epistemic uncertainty is the residual component isolated by subtracting the aleatoric uncertainty from the total predictive entropy. Hence, the calibration network of the GAN-based CAR model of Kassapis et al. (2021) (discussed in Section 4.2.1) is actually more inclined to model the epistemic rather than aleatoric uncertainty. This dependency is also demonstrated by the conflicting ideas on uncertainty encapsulated with TTA. In some works, it is argued that this approach models the aleatoric uncertainty in a better way (Wang et al., 2019a; Ayhan and Berens, 2022; Zhang et al., 2022a; Wang et al., 2019b; Rakic et al., 2024; Whitbread and Jenkinson, 2022; Roshanzamir et al., 2023). However, literature opposing this also exist and even suggests that TTA enables the modeling of epistemic uncertainty (Hu et al., 2022). Kahl et al. (2024) draw parallels with BNNs and demonstrate that
where $t$ are sampled augmentations from the input space $\mathcal{T}$. A well-trained model should be invariant to data augmentations. Consequently, predictive variance across augmentations is treated as irreducible aleatoric uncertainty, which is quantified by a non-zero expected entropy over these predictions. Non-zero mutual information indicates the model has not yet learned invariance to the augmentations and can be reduced with further training, deeming it to be epistemic. This decomposition is analogous to the one for conventional BNNs (Equation 3), while it contrasts with the decomposition used for generative models (Equation 41). The key message thus far is that, after considering the type of uncertainty to model, it is equally important to carefully choose the appropriate method for its representation, as confusion and misunderstandings can cause deviation from the targeted uncertainty.
Intertwinedness
@@
Despite the seemingly rigorous and well-defined distinction, it can be seen that implementation rapidly blurs this notion. A systematic study has found that the ability to separate the uncertainties can strongly depend on the data (Kahl et al., 2024). These findings are also supported by research presenting a strong correlation between epistemic and aleatoric uncertainty quantification methods (de Jong et al., 2024; Mucsányi et al., 2024), indicating that they represent uncertainty of similar kind. Kendall and Gal (2017) have also observed that explicitly modeling a single uncertainty tends to compensate for the lack of the other. In probabilistic models, we can also find that the parameter modeling methods, traditionally used for epistemic uncertainty, have been used to model observer variability, by for example using MC-Dropout or ensembling (see Table 1). Furthermore, using a combination of distribution and paramter modeling techniques can also beneficial (Gao et al., 2022). These findings further highlight the notion of the uncertainties being intertwined. On a slight side note, it is notable that popular BNN methods such as the Local Reparameterization trick use stochasticity in the output features of intermediate layers (usually considered to encapsulate aleatoric uncertainty) as a proxy to model the weight distributions (Kingma et al., 2015), further supporting this argument.
Modeling vs. quantifying
@@
A common misconception pertains that modeling a specific uncertainty also implies that the predicted output distribution represents that same uncertainty. This undermines the entropy decomposition in Equation (3) distinguishing between encapsulating, i.e. targeting an intended uncertainty type, versus quantifying, i.e. measuring the modeled uncertainty in isolation. For example, the estimation of mutual information is required to express the epistemic uncertainty in isolation. Nonetheless it is common for papers to quantify the uncertainty by inspecting the output distribution subject through repetitive sampling from the parameters densities, thereby approximating the predictive entropy encapsulating both uncertainties (Camarasa et al., 2021; Seeböck et al., 2019; Roshanzamir et al., 2023; Roy et al., 2018a; Rumberger et al., 2020b; Hoebel et al., 2022; Seeböck et al., 2019; Wundram et al., 2024b; Rumberger et al., 2020a; Sedai et al., 2018). This similarly holds for literature emphasizing aleatoric uncertainty when estimating the predictive variance (Zhang et al., 2022d; Monteiro et al., 2020; Valiuddin et al., 2024b; Gao et al., 2022). Since predictive variance can encompass both uncertainties, a crucial distinction must be made. Targeting a specific uncertainty for a task through either parameter or feature modeling is different from truly quantifying a type in isolation, which may require separate validation.